I need help with this word problem - A bank rober holds upa bank and tells the teller to hand over all her $100 bills and $500 bills. THe teller gave him 8 more $100 bills than $500 bills. If he gets away with $6,200, how many of each bill did the teller give him?
Answer by Earlsdon(6294)

Let x = the number of $100-bills and y = the number of $500-bills.
The problem states that:
x = y+8 "The teller gave him 8 more $100-bills than $500-bills" also...
x($100) + y($500) = $6200 Substitute x = y+8 and solve for y.
(y+8)($100)+y($500) = $6200 Simplify.
100y+800+500y = 6200
600y+800 = 6200 Subtract 800 from both sides.
600y = = 5400 Divide both sides by 600
y = 9 and x = y+8 = 17
The teller gives the bank robber 17 $100-bills and 9 $500-bills.
Check:
17($100)+9($500) = $1700+$4500 = $6200